OpenStudy (princeevee):

Find the quotient (9^2- 3b) ÷ b

OpenStudy (astrophysics):

\[\frac{ 9^2-3b }{ b } \implies \frac{ 81 }{ b }-\frac{ 3b }{ b }\]

OpenStudy (anonymous):

Give me a sec I'll work it out...

OpenStudy (princeevee):

Thanks

OpenStudy (anonymous):

\[\frac{ 81 }{ b }-3\] I believe that should be the right answer..

OpenStudy (princeevee):

thnk u so much

OpenStudy (anonymous):

No problem! :)

OpenStudy (astrophysics):

You do realize I did everything and you just had to cancel out the b, getting someone else to do your whole problem isn't going to help you.